OpenStudy (anonymous):

Which graph represents the solution set for the inequality 16 ≥ -3x + 4 > -17?

OpenStudy (texaschic101):

16 >= -3x + 4 16 - 4 >= -3x 12 >= -3x -4 <= x ( I switched the sign because I divided by a negative) -3x + 4 > - 17 -3x > -17 - 4 -3x > - 21 x < 7 so now we have x being less then 7 and greater or equal to -4 -4 <= x < 7 open circle on 7, closed circle on -4, shading in between

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

ty

OpenStudy (texaschic101):

no problem :)